OpenStudy (anonymous):

as the limit t approaches -2 at (t^3+8)/(t+2)

OpenStudy (anonymous):

factor the top and it looks like you might be able to cancel t+2 in the denominator

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

you can do it throw a chart if you want to do it that way

OpenStudy (anonymous):

is the limit 0?

OpenStudy (anonymous):

Just do polynomial division! t+2 is a factor of t^3+8 because -2 is a zero of t^3+8

OpenStudy (anonymous):

And then you just plug in t=-2 in the quadratic polynomial you get.

OpenStudy (anonymous):

I did, i got 0.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

is the answer.

OpenStudy (anonymous):

\[\lim_{t \rightarrow -2} t^2-2t+4\]

OpenStudy (anonymous):

btw correct term would be long division =P lol

OpenStudy (anonymous):

(x^3 +y^3)=(x+y) (x^2 -xy + y^2) t^3+8 is t^3 +2^3 (t+2)(t^2-2t+4) /(t+2) = (t^2-2t+4) provided t not equal to -2 as t gets close to -2 but is not equal to -2, the expression gets closer to 12.

OpenStudy (anonymous):

if you want to be real smart call it polynomial long division if you have to get it right -.-